Write the equation of the L satisfying the given geometric conditions.
L has y-intercept (0,-3) and is parallel to the line with equation y=2/3x+1

Answer by stanbon(75887)

L has y-intercept (0,-3) and is parallel to the line with equation y=2/3x+1
------------
If it is parallel, it has slope = (2/3)
You have the slope and the intercept
so, y = (2/3)x-3 is the equation you want.
===============
Cheers,
Stan H.